OpenStudy (anonymous):

(1+cos x+sin x)/ (1+cos x-sin x)= Sec x+Tan x

OpenStudy (cwrw238):

Right side = 1/cosx + sinx / cos x = ( 1 + sinx) / cos x) i'll write sin x as s and cos x as c so left side = (1 + c+ s) / (1 + c- s) = (1 + s) / c (right side)

OpenStudy (anonymous):

I got that for the right side but I can't figure out how to get the left to that point as well

OpenStudy (cwrw238):

now if we cross multiply c(1 + c+s) = (1 + s)(1 + c- s) c + c^2 + cs = 1 + c - s + s + sc - s^2 c + cs = 1 + c - s + s + sc - (s^2 + c^2) s^2 + c^2 = sin^2 x + cos^2 x = 1 so c + cs = 1 -1 - s+ s + c + cs c + cs = c + cs which completes the proof

OpenStudy (cwrw238):

you need to simplify and use the identity sin^x = cos^x = 1

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\dfrac{1+cosx+sinx}{1+cosx-sinx}\\~\\ =\dfrac{1+cosx+sinx}{1+cosx-sinx}\times \dfrac{secx}{secx}\\~\\ =\dfrac{secx+tanx+1}{secx-tanx+1}\\~\\ =\dfrac{secx+tanx+(sec^2x-tan^2x)}{secx-tanx+1}\\~\\ =\dfrac{(secx+tanx)(1+secx-tanx)}{secx-tanx+1}\\~\\ =(secx+tanx)\\~\\}\end{align}\)

OpenStudy (anonymous):

mathmath333 why would you multiply by secx/secx?

OpenStudy (mathmath333):

because that will give the next step \(\Large =\dfrac{secx+tanx+1}{secx-tanx+1}\\~\\\) which will be easy for ur proof

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{remember\\ \Large cosx\times secx=1\\~\\ \Large sinx\times secx=tanx}\end{align}\)